In what proportion should 5% and 1% hydrocortisone ointments be mixed to prepare a 2.5% ointment?
Answer #1

- By Allegation method & CA = 5.1. 72,5-1=1.5 Panty of A 82 - 3 lanky B F=2.5% = 11. 5-2.5-2.5 Pants of B*21 & Panty total =5